An astronaut working on the Moon tries to determine the gravitational constant G by throwing a Moon rock of mass m with a velocity of v vertically into the sky. The astronaut knows that the Moon has a density ρ of 3340 kg/m3 and a radius R of 1740 km.(a) Show with (1) that the potential energy of the rock at height h above the surface is given by:
4πG R3
E = −
3 mρ · R + h (2)
(b) Next, show that the gravitational constant can be determined by:
3 v2 .
R Σ−1
(c) What is the resulting G if the rock is thrown with 30 km/h and reaches 21.5 m?

1
Expert's answer
2021-03-04T15:18:51-0500

a) the gravitational potential energy at the distance r is


"E=-\frac{GMm}{r}=-\frac{\frac{4}{3}\pi R^3\rho Gm}{r}"

If r = R + h,


"E=-\frac{\frac{4}{3}\pi R^3\rho Gm}{R+h}"

b) At the surface of the Moon the potential energy is


"E=-\frac{\frac{4}{3}\pi R^3\rho Gm}{R+0}"

The total energy is


"K+E=0.5mv^2-\frac{4}{3}\pi R^2\rho Gm=-\frac{\frac{4}{3}\pi R^3\rho Gm}{R+h}"

"G=\frac{3v^2}{8\pi \rho R^2}\left(1-\frac{R}{R+h}\right)^{-1}"

c)


"G=\frac{3(\frac{30}{3.6})^2}{8\pi (3340)(1.74⋅10 ^
6)^2}\left(1-\frac{1.74⋅10 ^
6}{1.74⋅10 ^
6+21.5}\right)^{-1}\\=6.64\cdot10^{-11}\frac{Nm^2}{kg^2}"